OpenStudy (anonymous):

How many grams of NaCl will be produced when 2235 g of HCl are neutralized by an excess of NaOH according to the equation below? HCl + NaOH ---> H2O + NaCl

OpenStudy (anonymous):

36.5g HCl + 40gNaOH → 18g H2O + 58.5g NaCl. so it would be X / 58.5gNaCl = 2235gHCl / 36.5gHCl X = 58.5 x 2235 / 36.5 = Follow the same procedure for the remaining equations 2KNO3 = 2 x (39 + 14 + 3x16) = 202 2KNO2 = 2 x (39 + 14 + 2x16) = 170 X / 202gKNO3 = 536gKNO2 / 170gKNO2, solve for X I hope this helps!!!

OpenStudy (anonymous):

except you need to use the coefficient when calculating the mass in the balanced equation. Do you understand?

OpenStudy (anonymous):

Glad I could help